OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 + 10x2 + 24x ? –6, 0, 4 6, 0, –4 –6, 0, –4 6, 0, 4 Can someone please explain. would it be C or D?

OpenStudy (anonymous):

You can factor out an x and then have a quadratic to solve. Can you solve quadratic equations or do you need help?

OpenStudy (anonymous):

help pleaese

OpenStudy (anonymous):

x^3 + 10x^2 + 24x = 0 becomes x(x^2 + 10x + 24) = 0 which is considerably easier. We could do this because there was an "x" in each of the 3 terms. Also, it becomes obvious that x=0 is a solution. So you only have to find 2 more zeros. and you only have to look at x^2 + 10x + 24 = 0 which can be factored and takes the form of (x + a)(x + b) = 0. So, look for 2 numbers that when multiplied are 24 but when added are 10. That will give you "a" and "b". You then take the NEGATIVE of these values for the other 2 values that you need.